# Continual Learning from Simulated Interactions via Multitask Prospective Rehearsal for Bionic Limb Behavior Modeling

### v3

@@ A.4 Stability against forgetting
In addition, to systematically analyze the model performance as it is trained for multiple tasks, we also computed two metrics (similar to Díaz-Rodríguez et al. (2018)) to measure how well the model performs on previously trained tasks after training it with new tasks.
@@ B.1 Experimental setup
Besides the quantitative stability comparison of gait as reported in the main paper, we also analyzed data on trunk posture during two experimental scenarios: one with the subjects walking using the model-predicted ankle positions with the exoskeleton, and another where they walked without any assistance. We additionally looked at the variability in trunk kinematics, which is another critical indicator of balanced gait (Rábago et al., 2015). Figure 11 shows the lateral trunk angles in the case where a subject walked freely (blue) and when the subject walked with the assisted powered exoskeleton, whose ankle-joint positions were predicted from our model (red). Our findings revealed that the lateral trunk angles during assisted and unassisted walking were comparable, suggesting that the exoskeleton, guided by our model’s predictions, did not necessitate any compensatory movements from the subjects. This similarity implies that our model can effectively predict control signals that facilitate a stable and comfortable walking experience.
@@ C.1 Continual learning strategies
We compared our method to other continual learning strategies such as synaptic intelligence (SI) (Zenke et al., 2017), elastic weight consolidation (EWC) (Kirkpatrick et al., 2017), experience replay (ER), and gradient episodic memory (GEM)(Lopez-Paz & Ranzato, 2017).
@@ Appendix D Training and Hyperparameters
Rehearsal sampling strategy We employ a task-balanced reservoir sampling strategy for rehearsal, which ensures that samples from each task are maintained in a balanced manner over time. Specifically, this approach allows us to continuously update the rehearsal buffer with representative samples from all tasks, without favoring recent tasks over older ones. During each minibatch update, a balanced subset of samples from all tasks is used to ensure that the model’s performance does not degrade on earlier tasks. This method effectively mitigates the issue of bias towards more recent tasks. Moreover, our strategy aligns with well-established practices in continual learning, where task-balanced replay buffers are commonly used to prevent catastrophic forgetting Merlin et al. (2022); Krawczyk & Gepperth (2024).
@@ Appendix E Broader Impact
The strategies introduced in this paper have broad implications across human-centered assistive technologies, intelligent interfaces, and autonomous systems. Beyond prosthetics, this framework can improve the traversability of autonomous agents, such as legged robots Dey et al. (2022); Morrell et al. (2024) and self-driving vehicles Schmid et al. (2022), by enabling them to predict and adapt to dynamic environments. Furthermore, the ability to anticipate and account for the prediction errors is valuable for rehabilitation robotics and personalized healthcare solutions, offering more responsive and robust systems for patient care. Additionally, these modeling techniques may be adapted for industrial robotics to enhance safety in human-robot collaboration and for teleoperation scenarios that demand precise movement coordination. Taken together, this research represents a step toward resilient, human-centric AI technologies with wide-reaching societal benefits.

### v4

@@ A.4 Stability against forgetting
In addition, to systematically analyze the model performance as it is trained for multiple tasks, we also computed two metrics (similar to [52]) to measure how well the model performs on previously trained tasks after training it with new tasks.
@@ B.1 Experimental setup
Besides the quantitative stability comparison of gait as reported in the main paper, we also analyzed data on trunk posture during two experimental scenarios: one with the subjects walking using the model-predicted ankle positions with the exoskeleton, and another where they walked without any assistance. We additionally looked at the variability in trunk kinematics, which is another critical indicator of balanced gait [53]. Figure 11 shows the lateral trunk angles in the case where a subject walked freely (blue) and when the subject walked with the assisted powered exoskeleton, whose ankle-joint positions were predicted from our model (red). Our findings revealed that the lateral trunk angles during assisted and unassisted walking were comparable, suggesting that the exoskeleton, guided by our model’s predictions, did not necessitate any compensatory movements from the subjects. This similarity implies that our model can effectively predict control signals that facilitate a stable and comfortable walking experience.
@@ C.1 Continual learning strategies
We compared our method to other continual learning strategies such as synaptic intelligence (SI) [26], elastic weight consolidation (EWC) [25], experience replay (ER), and gradient episodic memory (GEM)[31].
@@ Appendix D Training and Hyperparameters
Rehearsal sampling strategy We employ a task-balanced reservoir sampling strategy for rehearsal, which ensures that samples from each task are maintained in a balanced manner over time. Specifically, this approach allows us to continuously update the rehearsal buffer with representative samples from all tasks, without favoring recent tasks over older ones. During each minibatch update, a balanced subset of samples from all tasks is used to ensure that the model’s performance does not degrade on earlier tasks. This method effectively mitigates the issue of bias towards more recent tasks. Moreover, our strategy aligns with well-established practices in continual learning, where task-balanced replay buffers are commonly used to prevent catastrophic forgetting [32; 54].
@@ Appendix E Broader Impact
The strategies introduced in this paper have broad implications across human-centered assistive technologies, intelligent interfaces, and autonomous systems. Beyond prosthetics, this framework can improve the traversability of autonomous agents, such as legged robots [55; 56] and self-driving vehicles [57], by enabling them to predict and adapt to dynamic environments. Furthermore, the ability to anticipate and account for the prediction errors is valuable for rehabilitation robotics and personalized healthcare solutions, offering more responsive and robust systems for patient care. Additionally, these modeling techniques may be adapted for industrial robotics to enhance safety in human-robot collaboration and for teleoperation scenarios that demand precise movement coordination. Taken together, this research represents a step toward resilient, human-centric AI technologies with wide-reaching societal benefits.